# Permutation twisted cohomology, remixed

### v2

@@ Abstract.
We generalize Balmer and Gallauer’s (permutation) twisted cohomology ring, working toward an alternative method of deducing the Balmer spectrum of the derived category of permutation modules for any finite $p$-group. The construction comes equipped with a canonical comparison map from the Balmer spectrum to the homogeneous spectrum of the twisted cohomology ring. We show the comparison map is injective for any finite $p$-group and furthermore, an open immersion when the twisted cohomology ring is noetherian. For elementary abelian $p$-groups, the twisted cohomology ring coincides with Balmer and Gallauer’s original construction.
To perform this construction, we utilize endotrivial complexes (i.e. the invertible objects of the derived category of permutation modules) arising up to a shift from Bredon homology of representation spheres. This topological structure allows us to construct certain $p$-local isomorphisms, from which we build a refined open cover of the Balmer spectrum indexed by conjugacy classes of subgroups of $G$. Under this open cover, every endotrivial is isomorphic to a shift of the tensor unit in each localization, thus verifying that all endotrivials are line bundles. When the twisted cohomology ring is noetherian, this open cover endows the Balmer spectrum with Dirac scheme structure.
@@
permutation module, twisted cohomology, comparison map, Balmer spectrum, tensor-triangular geometry
@@ Introduction
Let $k$ be a field of prime characteristic $p$ and $G$ be a finite group. Permutation $kG$-modules (modules which admit a $G$-stable $k$-basis), and their direct summands, $p$-permutation kGkG-modules, are perhaps the second easiest class of representations of $G$ over $k$ to define. In a wildly vast representation-theoretic cosmos of $kG$-modules, our $p$-permutation modules may seem tiny and insignificant, drowning in a sea of symmetries. However, our protagonists, or more precisely, the bounded homotopy category of $kG$-modules $\mathscr{K}(G):=\operatorname{K}_{b}(p\operatorname{-perm}(kG))$, exhibits significant control over the representation theory of $kG$ in general. For instance, if $G$ has an abelian Sylow $p$-subgroup $S$, Broué’s abelian defect group conjecture predicts a triangulated equivalence $\operatorname{K}_{b}(p\operatorname{-perm}(kGb_{0}))\cong\operatorname{K}_{b}(p\operatorname{-perm}(kN_{G}(S)c_{0}))$ where $b_{0}$ and $c_{0}$ denote the principal blocks of $kG$ and $kN_{G}(S)$ respectively. Perhaps more surprisingly, Balmer and Gallauer recently demonstrated that $\mathscr{K}(G)$ admits the bounded derived category $\operatorname{D}_{b}(\text{mod}(kG))$ as a Verdier quotient [BG23a], strengthening an unpublished result of Rouquier [Rou06]. Although these $p$-permutation are few, they are mighty; together, they recover the entire representation theory of $kG$.
If the reader is less representation-theoretically inclined, fear not! Our $p$-permutation category $\mathscr{K}(G)$ is the compact part of a “big” (i.e. rigidly-compactly generated) tt-category $\mathscr{D}(G):=\operatorname{DPerm}(G;k)$, the derived category of permutation modules. This category $\mathscr{D}(G)$ is equivalent to numerous big tt-categories arising from seemingly disparate areas of mathematics, such as:
@@
    the derived category of cohomological $k$-linear Mackey functors over $G$;
@@
    the homotopy category of modules over the constant Green functor $\operatorname{H}\underline{k}$ in genuine $G$-spectra;
@@
    the category of $k$-linear Artin motives generated by motives of intermediate fields in a Galois extension with Galois group $G$.
See [BG23b] for an overview of these correspondences; we remark Fuhrmann recently ascended equivalence (b) to the $\infty$-categorical level [Fuh25]. Hence any results concerning $\mathscr{D}(G)$ and $\mathscr{K}(G)$ translate over to the aforementioned categories and their compact parts respectively.
Continuing their foray into the land of permutation modules, Balmer-Gallauer deduced the Balmer spectrum $\operatorname{Spc}(\mathscr{K}(G))$ of $\mathscr{K}(G)$ in [BG25], a landmark achievement. Their construction of permutation twisted cohomology, a modification of the usual cohomology ring, plays a key role in this classification. The authors consider morphisms from the tensor unit to shifts of certain invertible objects (i.e. endotrivial complexes) corresponding to subgroups $N\triangleleft G$ of index $p$. The collection of these morphisms forms a multigraded ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, the twisted cohomology ring, extending the usual notion of a cohomology ring of a triangulated category, and comes with a canonical comparison map,
@@
which was previously constructed in [Bal10a]. A twist of the cohomology ring is necessary - since we are considering a homotopy category, the usual cohomology ring $\operatorname{End}^{\scriptscriptstyle\bullet}_{\mathscr{K}(G)}(k)$ is a tad drab on its own. The miracle of this construction is:
@@
[BG25, Theorem 10.5] Let $E$ be a finite elementary abelian $p$-group. The comparison map $\operatorname{comp}_{E}\colon\operatorname{Spc}(\mathscr{K}(E))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E))$ identifies $\operatorname{Spc}(\mathscr{K}(E))$ with an open subspace of the homogeneous spectrum of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E)$.
This remarkable fact, along with a colimit theorem [BG25, Theorem 11.10] reducing to the case of $G$ elementary abelian á la Quillen stratification, completes the description of $\operatorname{Spc}(\mathscr{K}(G))$ as a topological space, and as a Dirac scheme in the sense of [HP23]. As a set, $\operatorname{Spc}(\mathscr{K}(G))\cong\bigsqcup_{H\in s_{p}(G)/G}\operatorname{Spc}(\operatorname{D}_{b}(kG))$ ([BG25, Theorem 7.16]); we set $V_{G}:=\operatorname{Spc}(\operatorname{D}_{b}(kG))$. Hausmann and Schwede further considered the twisted cohomology ring for elementary abelian $2$-groups [HS25] (in the Mackey functor setting), which they entitled “Representation-graded Bredon homology.” The authors determined a minimal generating set and proved that for $p=2$, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E)$ is nilpotent-free.
Separately, the author deduced the Picard group of $\mathscr{K}(G)$, i.e. the group of endotrivials, in a sequence of papers [Mil24, Mil25b, Mil25a]. Significant parts of the classification were also completed by Bachmann in his dissertation [Bac16] in the context of Artin motives. When $G$ is a finite $p$-group, morally the endotrivials arise from Bredon homology of representation spheres, and therefore have additional topological structure. This was also observed by Yalçin in the context of $G$-Moore spaces [Yal17], where an analogous classification result was determined. Therefore, one obtains a canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$ associated to real representations modulo Adams operations. Interestingly, for elementary abelian $p$-groups, this basis corresponds exactly to the endotrivials used by Balmer-Gallauer in their twisted cohomology ring. The natural question for one to ask is:
Question.
Can we recover the Balmer spectrum $\operatorname{Spc}(\mathscr{K}(G))$ by further twisting (or if you will, re-mixing) $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ by the endotrivial complexes arising (up to shifts) from representation spheres?
We consider twisting by shifts of complexes coming from genuine $\mathbb{R}G$-modules as opposed to virtual representations, as if we were to twist by all endotrivials, the graded ring becomes unmanageably large (c.f. [BG25, Remark 12.24]). The analogous comparison to make here is that the group cohomology ring $\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)$ is noetherian and recovers the Balmer spectrum of the derived category $\operatorname{D}_{b}(kG)$, whereas the Tate cohomology ring is too large. We call such endotrivials effective, and their corresponding h-mark homomorphisms are monotonically decreasing superclass functions. Similarly, Hausmann-Schwede call the corresponding “quadrant” of twisted cohomology associated to twists by effective endotrivials the effective cone.
Denote the subset of the canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$ arising from nontrivial irreducible real representations of $G$ by $\mathcal{B}(G)$. Let $\mathbb{N}^{\mathcal{B}(G)}=\{q\colon\mathcal{B}(G)\to\mathbb{N}\}$ be the monoid of twists, i.e. tuples of non-negative integers indexed by the set $\mathcal{B}(G)$. The monoid of twists is equivalently nothing more than the submonoid of $\operatorname{Pic}(\mathscr{K}(G))$ generated by $\mathcal{B}(G)$. Then the $(\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)})$-graded ring
@@
is the (re-)twisted cohomology ring; see Definition 5.1. We note that only non-positive shifts $s\leq 0$ play a nontrivial role.
We take the first steps towards answering this question for finite $p$-groups by proving injectivity of the comparison map. The $p$-group case is perhaps the most fundamental: given a finite group $G$ with $p$-Sylow $S$, one can determine $\operatorname{Spc}(\mathscr{K}(G))$ from $\operatorname{Spc}(\mathscr{K}(S))$ via restriction.
Theorem A. (Theorem 6.1).
Let $G$ be a finite $p$-group. The comparison map $\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ is injective.
Generally, injectivity of the comparison map can be a harder condition to prove (if it holds at all, which it frequently may not!) than surjectivity (after a categorical localization), which general tt-geometry usually resolves. Circling back to Balmer-Gallauer’s original construction, another feature of $\operatorname{Spc}(\mathscr{K}(E))$ for $E$ elementary abelian is a spectral open cover $\{U(H)\}_{H\leq E}$ of $\operatorname{Spc}(\mathscr{K}(E))$ indexed by subgroups of $E$. This cover has two remarkable features:
@@
    Over each open $U(H)$, all endotrivials for $E$ are trivial, i.e. isomorphic to a shift of the tensor unit in the localization $\mathscr{L}(H):=\mathscr{K}(E)|_{U(H)}$;

### v3

@@ Abstract.
For each endotrivial complex arising from Bredon homology of a representation sphere, we construct $p$-local quasi-isomorphisms, called forerunners, enabling us to extend Balmer–Gallauer’s results in [BG25, Part II] concerning the tensor-triangular geometry of permutation modules for elementary abelian $p$-groups to all $p$-groups. We construct an open cover of the Balmer spectrum under which all endotrivials are line bundles, that is, every endotrivial is locally isomorphic to a shifted tensor unit. We define a remixed permutation twisted cohomology ring for which the canonical comparison map from the Balmer spectrum to the homogeneous spectrum of the twisted cohomology ring is injective. If the twisted cohomology ring is Noetherian, the comparison map is an open immersion, and the open cover endows the Balmer spectrum with Dirac scheme structure. We prove Noetherianity holds for Dedekind groups and all $p$-groups of order at most $p^{3}$, and conjecture Noetherianity holds for all finite $p$-groups.
@@
permutation module, twisted cohomology, comparison map, Balmer spectrum, tensor-triangular geometry, endotrivial
@@
For the duration of this paper, $k$ denotes a field of prime characteristic $p$ and $G$ denotes a finite group.
@@ Introduction
It is oft claimed that permutation $kG$-modules, modules which admit a $G$-stable $k$-basis, and their direct summands, $p$-permutation kGkG-modules, are at minimum the second easiest class of representations of $G$ over $k$ to define. In a wildly (c.f. [BD77]) vast representation-theoretic cosmos of $kG$-modules, our $p$-permutation modules appear tiny and insignificant, drowning in the endless sea of symmetries. However, our protagonist, the bounded homotopy category of $kG$-modules $\mathscr{K}(G):=\operatorname{K}_{b}(p\operatorname{-perm}(kG))$, exhibits significant control over said cosmos. For a classical example, if $G$ has an abelian Sylow $p$-subgroup $S$, Broué’s abelian defect group conjecture [BRO90, RIC96] predicts a particular triangulated equivalence $\operatorname{K}_{b}(p\operatorname{-perm}(kGb_{0}))\cong\operatorname{K}_{b}(p\operatorname{-perm}(kN_{G}(S)c_{0}))$ where $b_{0}$ and $c_{0}$ denote the principal blocks of $kG$ and $kN_{G}(S)$ respectively; such an equivalence induces a corresponding derived equivalence. More recently, Balmer–Gallauer demonstrated that every $kG$-module has a finite resolution by $p$-permutation modules [BG23a]. Consequently, $\mathscr{K}(G)$ admits the bounded derived category $\operatorname{D}_{b}(\text{mod}(kG))$ as a Verdier quotient, strengthening an unpublished result of Rouquier [ROU06]. Although these $p$-permutation are few, they are mighty; together, they recover the entire representation theory of kGkG!
In a landmark series of papers [BG22, BG23a, BG23b, BG25], Balmer–Gallauer elucidated the tensor-triangular geometry of $\mathscr{K}(G)$. In particular, in [BG25], the authors deduced its Balmer spectrum $\operatorname{Spc}(\mathscr{K}(G))$, the universal support of the category. [^1] This classification has ramifications beyond representation theory: to the geometrically or topologically inclined reader, this classification can be interpreted in terms of cohomological Mackey functors, Artin motives, or the homotopy category of modules over the constant Green functor, see [BG23b, FUH25] As a set, the Balmer spectrum is built from $p$-local cohomology:
$\operatorname{Spc}(\mathscr{K}(G))=\bigsqcup_{H\in s_{p}(G)/G}\operatorname{Spc}(\operatorname{D}_{b}(k[{G}/\!\!/{H}]))\cong\bigsqcup_{H\in s_{p}(G)/G}\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H};k)).$
This stratification is obtained from certain tt-functors classically known as Brauer quotients for modules, and in [BG25] are termed modular fixed points, functors $\Psi^{H}\colon\mathscr{K}(G)\to\mathscr{K}({G}/\!\!/{H})$, where ${G}/\!\!/{H}:=N_{G}(H)/H$ is the Weyl group of $G$ at $H$. These functors mimic the fixed-point operation for $G$-sets: given any $G$-set $X$ and $p$-subgroup $H\leq G$, one has a natural isomorphism $\Psi^{H}(kX)\cong k[X^{H}]$. We’ll discuss the nitty-gritty of the classification, but first take a narrative detour to the domain of endotrivials.
Endotrivial complexes and forerunners
Separately, in a sequence of papers [MIL24, MIL25a, MIL25b], we deduced the Picard group $\operatorname{Pic}(\mathscr{K}(G))$, i.e., the group of endotrivial complexes for all finite groups. Endotrivials are by definition the invertible objects of $\mathscr{K}(G)$, the chain complexes $C\in\mathscr{K}(G)$ satisfying $C^{*}\otimes C\simeq k[0]$. This classification was also partially determined in Bachmann’s dissertation [BAC16] in the context of Artin motives. Independently, a similar computation was performed by Yalçin [YAL17] for $p$-groups in the context of $G$-Moore spaces, which was partially extended to finite groups in joint work with Gelvin [GY21].
There are multiple alternative ways to think about endotrivial that lead to a classification. In [MIL25a, MIL25b], we deduced $\operatorname{Pic}(\mathscr{K}(G))$ by considering endotrivials as a special class of more generally “relatively” endotrivial complexes akin to endopermutation modules (see [BOU10, Chapter 12]); this parallel to endopermutation modules is developed further in [MIL26]. Another characterization considered by Bachmann and Gelvin–Yalçin is topological in nature: endotrivials arise from Bredon homology of (virtual) representation spheres, i.e., one-point compactifications of real representations. In other words, one has a homomorphism $\operatorname{RO}(G)\to\operatorname{Pic}(\mathscr{K}(G))$. In fact, for $p$-groups (but not in general), the homomorphism is surjective, and its kernel is characterized by the Adams operation or the $j$-homomorphism. In particular, for $G$ a $p$-group, $\operatorname{Pic}(\mathscr{K}(G))$ has a canonical $\mathbb{Z}$-basis associated to the irreducible real representations of $G$.
In this paper, we adopt this topological point of view for $p$-groups, with a particular emphasis on effective endotrivials, the endotrivials arising from Bredon homology of (genuine) representation spheres. Such endotrivials have additional structure: they are chain complexes of free modules over the orbit category $\Gamma_{G}$. This structure allows us to construct certain $p$-local quasi-isomorphisms which we term forerunners, as they foretell quasi-isomorphisms prior to applying modular fixed points.
Theorem A.
Let $G$ be a finite $p$-group. For every effective endotrivial complex (see Definition 2.9) and subgroup $H\leq G$, there exists a forerunner homomorphism $\iota^{H}_{C}:k\to C[h_{C}(H)]$ such that $\Psi^{H}(\iota^{H}_{C})$ is a quasi-isomorphism (Theorem 4.3). Here, $h_{C}$ denotes the h-mark homomorphism of $C$, see Definition 2.1.
Let us justify to the inquisitive reader why the existence of forerunners is noteworthy. First, the existence of such morphisms is by no means a given: it is a special property of effective endotrivials. Proving their existence requires developing the theory of chain complexes of permutation $kG$-modules arising from the orbit category $\Gamma_{G}$; we do not list the results here, but believe they will be of independent interest. Second, Gallauer has already utilized these morphisms in his recent initiation of the study of periodicity for symmetric monoidal $\infty$-categories [GAL25]; we remark on this work in the sequel. But the real justification, the true power of these morphisms, is that they enable us to explain the geometry of permutation modules in a much more general light, as we’ll now explain.
A new open cover of the spectrum
In [BG25], Balmer–Gallauer split their deduction of $\operatorname{Spc}(\mathscr{K})$ into two parts: Part I concerns deducing $\operatorname{Spc}(\mathscr{K}(G))$ as a set and constructing the modular fixed-point functors, and Part II deduces the topology of $\operatorname{Spc}(\mathscr{K}(G))$. In Part II, the authors prove a colimit theorem a la Quillen stratification reducing classification to the case of $G$ elementary abelian, then investigate this case in greater depth. Using the classification of endotrivials, the existence of forerunners, and the topological origins of effective endotrivials, we extend these investigations in the context of any finite $p$-group. Our hope is that $\operatorname{Spc}(\mathscr{K}(G))$ can be completely deduced without any gluing. The $p$-group case is the crux: given a finite group $G$ with $p$-Sylow $S$, one can determine $\operatorname{Spc}(\mathscr{K}(G))$ from $\operatorname{Spc}(\mathscr{K}(S))$ by restriction via [BG25, Propositions 3.8, 4.7].
We first review Balmer–Gallauer’s methods and results. Let $E$ be an elementary abelian $p$-group. Balmer–Gallauer constructed an open cover $\{U(H)\}_{H\leq E}$ of $\operatorname{Spc}(\mathscr{K}(E))$ indexed by subgroups of $E$. Explicitly, these opens are obtained by considering the open loci of certain morphisms from the tensor unit to certain endotrivials. This cover has some remarkable features:
@@
    All endotrivials for $E$ are isomorphic to a shift of the tensor unit in the localization $\mathscr{K}(E)|_{U(H)}$; in other words, all endotrivials for are line bundles [BG25, Remark 13.10];
@@
    Each open $U(H)$ contains the closed point $\mathscr{M}(H)\in\operatorname{Spc}(\mathscr{K}(E))$. Hence $\{U(H)\}_{H\leq E}$ is an open cover of $\operatorname{Spc}(\mathscr{K}(E))$; [BG25, Proposition 13.11]
@@
    Under this open cover, there is an identification $\mathscr{K}(E)|_{U(1)}\cong\operatorname{D}_{b}(kE)$. In other words, $U(1)$ is nothing more than the cohomological open, corresponding to the image under $\operatorname{Spc}$ of the localization $\mathscr{K}(E)\twoheadrightarrow\operatorname{D}_{b}(kE)$; [BG25, Proposition 13.14]
This open cover is later used for to endow $\operatorname{Spc}(\mathscr{K}(E))$ with Dirac scheme structure in the sense of [HP23]; a Dirac scheme is to a usual scheme as a graded ring is to a non-graded one. The morphisms used to construct these opens are examples of forerunners - this observation suggests that such an open cover in fact exists in greater generality.
Theorem B.
Let $G$ be a finite $p$-group. There exists an open cover $\{U(H)\}_{H\in s_{p}(G)/G}$ (4.6) indexed by conjugacy classes of subgroups of $G$ for which the following holds.
  1. (a)
    Every endotrivial $C$ is a line bundle, i.e., over each open $U(H)$ is isomorphic to a shift of the tensor unit in the localization $\mathscr{K}(G)|_{U(H)}$. Explicitly, $C\cong k[h_{C}(H)]$ in $\mathscr{K}(G)|_{U(H)}$, where $h_{C}$ denotes the h-mark homomorphism associated to $C$; (Theorem 4.13)
  2. (b)
    Each open $U(H)$ contains a unique closed point $\mathscr{M}(H)\in\operatorname{Spc}(\mathscr{K}(G))$, corresponding to the unique closed point of $\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H};k))$; (Theorem 4.12)
  3. (c)
    The open $U(1)$ is the cohomological open $\operatorname{Spc}(\operatorname{D}_{b}(kG))\subseteq\operatorname{Spc}(\mathscr{K}(G))$. Therefore, we have a tt-equivalence $\mathscr{K}(G)|_{U(1)}\cong\operatorname{D}_{b}(kG)$. (Theorem 4.19)
(Remixed) permutation twisted cohomology and Dirac geometry
To completely deduce the topology of $\operatorname{Spc}(\mathscr{K}(E))$, Balmer–Gallauer define the permutation twisted cohomology ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. Explicitly, the authors consider morphisms from the tensor unit to shifts of certain endotrivials corresponding to subgroups $N\triangleleft G$ of index $p$. The collection of these morphisms defines the multigraded ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, and extends the usual notion of the cohomology ring of a triangulated category. The ring comes with a canonical comparison map,
@@
extending the usual comparison map of [BAL10a]. A twisted variant of the cohomology ring is necessary - since $\mathscr{K}(G)$ is a homotopy category of chain complexes, the usual cohomology ring $\operatorname{End}^{\scriptscriptstyle\bullet}_{\mathscr{K}(G)}(k)$ is a tad drab on its own. The miracle of this construction is:
@@
[BG25, Theorem 10.5] Let $E$ be an elementary abelian $p$-group. The comparison map $\operatorname{comp}_{E}\colon\operatorname{Spc}(\mathscr{K}(E))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E))$ identifies $\operatorname{Spc}(\mathscr{K}(E))$ with an open subspace of the homogeneous spectrum of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E)$. In particular, $\operatorname{Spc}(\mathscr{K}(E))$ is a Dirac scheme.
Hausmann–Schwede have further considered the twisted cohomology ring for elementary abelian $2$-groups in the context of Mackey functors [HS25]. The authors determined a minimal generating set and proved that for $p=2$, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E)$ is nilpotent-free.
Interestingly, for elementary abelian $p$-groups, the endotrivials involved in the construction of twisted cohomology are precisely the effective endotrivials we’ve considered, motivating our following “remixed” construction.
Definition (Definition 5.1).
Let $G$ be a $p$-group. Denote the subset of the canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$ arising from nontrivial irreducible real representations of $G$ by $\mathcal{B}(G)$. Let $\mathbb{N}^{\mathcal{B}(G)}=\{q\colon\mathcal{B}(G)\to\mathbb{N}\}$ be the monoid of twists, i.e., tuples of non-negative integers indexed by the set $\mathcal{B}(G)$. Then the $(\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)})$-graded ring
@@
is the (remixed) twisted cohomology ring.
We twist only by shifts of effective endotrivials rather than all endotrivials, as otherwise, the graded ring becomes unmanageably large (c.f. [BG25, Remark 12.24]). Analogously, the group cohomology ring $\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)$ is Noetherian and recovers the Balmer spectrum of the derived category $\operatorname{D}_{b}(kG)$, whereas the Tate cohomology ring $\hat{\operatorname{H}}{}^{\scriptscriptstyle\bullet}(G;k)$ is, for sufficiently interesting $G$, non-Noetherian and contains only nilpotents in negative degrees [BC92].
Theorem C.
Let $G$ be a finite $p$-group. The comparison map
$\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)),\,\mathscr{P}\mapsto\left\langle f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\text{ homog.}\mid\operatorname{cone}(f)\not\in\mathscr{P}\right\rangle$
is injective (Theorem 6.2) with dense image (Corollary 5.13). If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian, the comparison map is an open immersion. (Corollary 6.6)
The proof of this theorem is highly technical and relies on our forerunner homomorphisms and an induction argument up a subnormal series of $p$-subgroups. It is seemingly necessary to twist by all effective endotrivials (or at minimum, a full-rank subset) to obtain these properties; see for instance Remark 4.16. In this sense, we believe our construction for $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is the “correct” generalization.
Since every endotrivial is a line bundle with respect to our open cover $\{U(H)\}$ of Theorem B, for every open $U(H)$, the twisted cohomology ring locally identifies with the usual cohomology ring in the localization $\mathscr{K}(G)|{U(H)}$. This allows us to use the classical comparison map of [BAL10a] locally in each open. Morally, twisted cohomology globally describes the local data in each open. If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian, then we can endow $\operatorname{Spc}(\mathscr{K}(G))$ with Dirac scheme structure, as was the case for elementary abelian $p$-groups.
Theorem D.
Let $G$ be a finite $p$-group. If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian, the comparison map is an open immersion and for each subgroup $H\leq G$, restricts to a local homeomorphism
$\operatorname{comp}_{\mathcal{L}_{G}(H)}\colon U(H)\to\operatorname{Spec}^{h}(\operatorname{End}_{\mathscr{L}_{G}(H)}^{\scriptscriptstyle\bullet}(k)).$
In this case $(\operatorname{Spc}(\mathscr{K}(G)),\mathcal{O}_{G}^{\scriptscriptstyle\bullet})$ is a Dirac scheme. (Corollary 6.6, Corollary 6.7) Noetherianity holds for the following groups:
@@
    Res-onto $p$-groups (see Definition 7.1). In particular, all Dedekind $p$-groups, groups with every subgroup normal; (Theorem 7.4)